prove that every circle passing through points -1 and +1 have an equation of the form zz'+ikz-ikz'-1=0
where k is real and z' = z conjugate

Grade:

1 Answers

AskiitiansExpert Mohit-IITD
13 years ago

Dear Rachneet,

There are two approaches to this problem:

Approach 1: Using co-ordinate geomety

Family of circles using diametric form with points (-1,0) and (1,0) => x2+y2=1as diameter and line passing through these two points =>y=0 will be

x2+y2-1 +py=0 where p is a variable

=> |z|2-1+p(z-z')/2i=0

=> zz'+ikz-ikz'-1=0 where k=-p/2 is real

Approach 2:

Equation of circle: Arg[(z-1)/(z+1)]=θ

=> (z-1)/(z'-1)=[(z+1)/(z'+1)]ei2θ

=> Simplify to get (zz'-1)+(z'-z)[(ei2θ+1)/(ei2θ-1)]=0

=> (zz'-1)+(z'-z)(-icotθ)=0 where (ei2θ+1)/(ei2θ-1)=-icotθ

=> zz'+ikz-ikz'-1=0 where k=cotθ is real

 

Hope this helps.

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation. 

 

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.